OpenStudy (anonymous):

find f^-1 and verify that (fof^-1)(x)=(f^-1of)(x)=x f(x)=2x+3

OpenStudy (anonymous):

how would you solve \[2x+3=7\]?

OpenStudy (anonymous):

1) subtract 3 2) divide by 2 that is what the inverse does, subtract 3, divide by 2, i.e. \[f^{-1}(x)=\frac{x-3}{2}\]

OpenStudy (anonymous):

if you like you can write \[x=2y+3\] and solve for \(y\) you get \[x=2y+3\\x-3=2y\\ \frac{x-3}{2}=y\] i.e. \[f^{-1}(x)=\frac{x-3}{2}\]

OpenStudy (anonymous):

ok thank you. when the problem asks to find (fof^-1)(x)=(f^-1of)(x)=x, I know I am supposed to plug in the inverse in every x of the original equation and vice versa, but I'm not exactly sure how both will end up equaling just x.

OpenStudy (anonymous):

bunch 'o algebra

OpenStudy (anonymous):

For the verification part, you have \[f(x)=2x+3\\ f^{-1}(x)=\frac{x-3}{2}\] So, \[\begin{align*}\left(f\circ f^{-1}\right)(x)&=f\left(f^{-1}(x)\right)\\ &=f\left(\frac{x-3}{2}\right)\\ &=2\left(\frac{x-3}{2}\right)+3\\ &=x-3+3\\ &=x \end{align*}\] The procedure is identical for determining \(\left(f^{-1}\circ f\right)(x)=x\).

OpenStudy (anonymous):

\[f(f^{-1}(x))=2(\frac{x-3}{2})+3\]

OpenStudy (anonymous):

if your inverse is correct you will end up with a raft of cancellation, giving you nothing but \(x\) at the end

OpenStudy (anonymous):

alright thank you both so much. I was just really confused with the way the problem was written.